No idea how to do this can someone help me with thNo idea how to do this can someone help me with these 3 questions?
Answer #1

torque about point A = mg * perpendicular distance of the point from the force

torque about point A = mg * L / 2

torque about point A = mgL / 2 Nm

torque about point B = mg * perpendicular distance of the point from the force

torque about point B = mg * 0

torque about point B = 0 Nm

torque about point C = mg * perpendicular distance of the point from the force

torque about point C = mg * L * sin(120) / 2

torque about point C = 0.433 * mgL Nm